Please help me solve this equation:-2x^2-16x-10=20
Answer by helpnalgebra(91)

-2x^(2)-16x-10=20

-2x^(2)-16x-30=0
2x^(2)+16x+30=0
2(x^(2))+2(8x)+2(15)=0
2(x^(2)+8x+15)=0
2(x+5)(x+3)=0
(x+5)(x+3)=0
x+5=0
x=-5
x+3=0
x=-3
x=-5,-3 answer